A plane electromagnetic wave traveling in the positive  direction of an x axis in vacuum has components Ex = Ey = 0  and

Ez = (4.8 V/m) cos[(?
Answer #1
according to given problem , we have
x, y , z - components of the electeomagnetic waves is
Ex = Ey = 0 and Ez = (4.8 V/m) cos[(?*1015 s-1)(t - x/c)]
here , maximum electric field Em = 5 V/m
a)
the amplitude of the magnetic field is
Bm = Em/c
= (4.8 V/m) / ( 2.998*108 m/s)
= 1.601*10-8 T
b)
from given problem ,
the waves oscillates in z-direction and its travels in +ve x-direction.
so , Bx = Bz = 0
therefore , the oscillation of the magnetic field is parallel to the y-axis.
c)
electromagnetic wave propagation is determined by E*B.
( * = cross product)
if the electric field component is in the positive direction of the z axis ,
then the magnetic field is in the y-axis (direction)
By = Bm cos[(?*1015 s-1)(t - x/c)]
= (1.601*10-8 T) cos[(?*1015 s-1)(t - x/c)]
Answer #2
according to given problem , we have
x, y , z - components of the electeomagnetic waves is
Ex = Ey = 0 and Ez = (4.8 V/m) cos[(?*1015 s-1)(t - x/c)]
here , maximum electric field Em = 5 V/m
a)
the amplitude of the magnetic field is
Bm = Em/c
= (4.8 V/m) / ( 2.998*108 m/s)
= 1.601*10-8 T
b)
from given problem ,
the waves oscillates in z-direction and its travels in +ve x-direction.
so , Bx = Bz = 0
therefore , the oscillation of the magnetic field is parallel to the y-axis.
c)
electromagnetic wave propagation is determined by E*B.
( * = cross product)
if the electric field component is in the positive direction of the z axis ,
then the magnetic field is in the y-axis (direction)
By = Bm cos[(?*1015 s-1)(t - x/c)]
= (1.601*10-8 T) cos[(?*1015 s-1)(t - x/c)]